OpenStudy (anonymous):

Elsa runs 3 miles in 25 minutes. At the same rate, how many miles would she run in 20 minutes?

OpenStudy (anonymous):

Do you have any idea how to start?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Okay well there's two ways to do this, but I'm pretty sure you're doing cross multiplying.

OpenStudy (anonymous):

What you can do is set up an equation

OpenStudy (anonymous):

You know she ran 3 miles in 25 minutes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 3 }{ 25 }\]

OpenStudy (anonymous):

8.3?

OpenStudy (anonymous):

Okay so \[\frac{ 3 }{ 25 }= \frac{ ? }{ ? }\]

OpenStudy (anonymous):

every mile?

OpenStudy (anonymous):

That was the other way to do it lol

OpenStudy (anonymous):

but you're most likely doing it this way

OpenStudy (anonymous):

So let me finish with this way and then I'll explain the second way

OpenStudy (anonymous):

sounds good '

OpenStudy (anonymous):

Okay so we have the first equation

OpenStudy (anonymous):

Do you know what goes into the second equation?

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

Okay I'll help you out with that

OpenStudy (anonymous):

\[\frac{ 3 }{ 25 }= \frac{ x }{ 20}\]

OpenStudy (anonymous):

You know she's running 20 minutes so you know the bottom is 20

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Now cross multiply

OpenStudy (anonymous):

2.4?

OpenStudy (anonymous):

You got it!

syed98 (syedmohammed98):

use the ratio method 

OpenStudy (anonymous):

so 2.4 miles in 20 mins right?

OpenStudy (anonymous):

The other way was to just divide 3/25 and which would you get 0.12 which is what she runs per minute and then multiply that by 20

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problemo! It's definitely 2.4 I checked myself :)

syed98 (syedmohammed98):

Your'e Welcome :)

OpenStudy (k_lynn):

Nice explanations, guys :D

syed98 (syedmohammed98):

Thanks @k_lynn